OpenStudy (anonymous):

integral of = t^2 square root of t^3+1 dt .?

OpenStudy (mimi_x3):

\[\large \int\limits\frac{t^{2}}{\sqrt{t^{3}+1}} dx\]?

OpenStudy (lgbasallote):

\[\huge \int t^2 \sqrt{t^3+1}dt\] let u = t^3 + 1 du = 3t^2dt reqrite... \[\LARGE \frac13 \int \sqrt u du\] got it?

OpenStudy (mimi_x3):

lol woops im blind

OpenStudy (anonymous):

nope the t^2 is not over the the square root of t^3+1. :)

OpenStudy (lgbasallote):

do you get my solution the @Beatles ?

OpenStudy (anonymous):

yup but what is the answer? :)

OpenStudy (anonymous):

please subtitue the value. :)

OpenStudy (lgbasallote):

what do you mean?

OpenStudy (mimi_x3):

well tell us what you got for the answer (:

OpenStudy (anonymous):

Never mind. The answer is 1/3 (t^3+1)^3/2 :) is it correct?

OpenStudy (lgbasallote):

i dont think so..

OpenStudy (anonymous):

opps. I'm wrong. :) this 2/9 (t^3+1)^3/2

OpenStudy (lgbasallote):

+C haha

OpenStudy (lgbasallote):

and yeah that's right

OpenStudy (anonymous):

haha yah. :)